OpenStudy (anonymous):

A popular newspaper puzzle involves a series of letters that can be rearranged to form a word. Will is writing his own version of the game for a school project. He wants to scramble the following word for his puzzle. NUMBERS How many ways are there to arrange the letters with the letter b as the first letter? Work please!

OpenStudy (anonymous):

bum bun buns bums

OpenStudy (anonymous):

..

OpenStudy (anonymous):

well, most of these are going to be gibberish letters, but starting with B the total number of combinations you can make are.... look, so B, and then theres 6 letters left, and so you take the 6 letters and put them to the power of 6

OpenStudy (anonymous):

so it would me 6 to the power of 6?

OpenStudy (anonymous):

Um, I think so, yes...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

like the blind man and the elephant, everyone interprets this problem differently what is it really asking? how many ways to scramble the letters in the word NUMBERS if the first letter is B?

OpenStudy (valpey):

After the B, there are 6 distinct letters with 6! arrangements. 6! = 720

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

if that is the question, then the answer is what valpey said

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

as in \[6\times 5\times 4\times 3\times 2\]

OpenStudy (anonymous):

Oh, I see.